OpenStudy (anonymous):

144 employees is 9% of ____ employees

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (haseeb96):

just multiply 144 with 0.09 and get your answer

OpenStudy (anonymous):

fr?

OpenStudy (haseeb96):

fr ??? what?

OpenStudy (anonymous):

\[144=0.09\] \[\frac{ 144 }{ 9 }=0.01\] then multiply by 100

OpenStudy (anonymous):

1600?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay can you help me with another?

OpenStudy (anonymous):

ok

OpenStudy (haseeb96):

oh sorry i made mistake in reading question

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

During a sale, a dress decreased in price from $80 to 72. what percent did it decrease?

OpenStudy (anonymous):

\[(1-\frac{ 72 }{ 80 }) \times 100\]

OpenStudy (anonymous):

is it 10?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw